what is the molality of 40.0g KBr dissolved in 800.0g water

1
Expert's answer
2022-03-31T11:22:04-0400

Solution:

Molality = Moles of solute / Kilograms of solvent


solute = KBr

solvent = H2O (water)


The molar mass of KBr is 119 g/mol

Therefore,

Moles of KBr = (40.0 g KBr) × (1 mol KBr / 119 g KBr) = 0.336 mol KBr

Moles of KBr = 0.336 mol


Kilograms of water = (800.0 g H2O) × (1 kg / 1000 g) = 0.8 kg H2O


Thus:

Molality of KBr solution = Moles of KBr / Kilograms of water

Molality of KBr solution = (0.336 mol) / (0.8 kg) = 0.42 mol/kg = 0.42 m

Molality of KBr solution = 0.42 m


Answer: The molality of KBr solution is 0.42 m